# Tunneling method for Hawking quanta in analogue gravity

### v1

@@ VI Particle creation: subcritical flow
Now, we complete our analysis by investigating a subcritical flow for which $|v|<1$ besides $|v_{\mathrm{max}}|\sim O(1)$. Then, no Killing/sonic horizon exists, hence such a flow mimics a very compact star rather than a black hole. This is a quite interesting case, given that experimental problems in shallow water waves experiments concerning the stability of horizons led to the prevalent realisation of such “near-critical” flows  [37, 32, 38, 39]. Experiments that nonetheless observed particle production. This led to several numerical [18, 19] and analytical investigations [31] which verified the presence of a particle flux albeit generically characterised by substantial deviations from thermality.
Here we shall re-analyse this phenomenon using the tunneling method with the above introduced approximant trajectory. Doing so we shall recover the qualitative behaviour of the aforementioned results and at the same time clarify the basic physics behind this “at first sight puzzling” particle production. Let us start with the EFH condition (25).
@@
Before particularising to a specific profile of $v(x)$, we quantify first how far from the sonic point $v=-1$ can the flow most negative value be so to still admit some solution to (25) for our subluminal dispersion relation. Within our setting this yields numerically the upper bound (namely, the solution of (25) for $\alpha=1/2$)
@@
In order to proceed further in our investigation, let us now assume a velocity profile of the form
@@
Figure 9: $\kappa(\alpha)$ for different values of $\varepsilon$: increasing $\varepsilon$ reduces the range of $\alpha$ for which (25) has a solution. Moreover, $\kappa(\alpha)$ starts from 00 with a nonvanishing derivative (apart from the critical case $\varepsilon=0$) which implies a deviation from thermality in the emission that will be the topic of the last section
@@ VII Phenomenological Considerations
In this section, we highlight some important aspects and subtelties related to our approach to provide a complete and comprehensive picture. For instance, we analyse the properties of the spectrum and the internal consistency of the methodology.
@@ VII.3 Energy conservation: subcritical flow-subluminal dispersion relation
  • the lower branch of the negative energy mode for the subluminal case (which is the lingering mode in Fig.1) is dispersive around $v=0$ (the origin of the axes in Fig.1). In (18) we have shown that the regular solutions around $v\simeq 0$ consist in two positive-$\Omega$ solutions, one ingoing and one outgoing. However, none of the lower branches of the lingering or the turning mode reach $v=0$, meaning that those branches are nonregular in that region of spacetime. This can be explained by the following fact: for these two branches $|k|$ increases while approaching $v=0$, so much so that for the subluminal dispersion relation $\omega(k)$ becomes imaginary and no real mode of these branches can be further propagated towards $v=0$.
@@ VIII Discussion
This article should serve as a primer for Hawking radiation in acoustic gravity, that provides an accessible and physically intuitive description via the tunneling method. Due to its simplicity and versatility, this formalism allowed us to explore several different scenarios that may be prevalent in analogue gravity set-ups. We investigated three different flow types — supercritical, critical, and subcritical — over which we analysed the particle creation for a scalar field with generalised dispersion relations of subluminal as well as superluminal type. As a result, we found that particle creation can be associated with an effective horizon experienced by a trajectory that interpolates those of the Hawking partners outside and inside the (effective) horizon. This approximant only fails to track these particles in a region of size below the de Broglie wavelength (where they cannot be treated as separated on-shell particles). The aforementioned effective horizon is always located inside the Killing horizon for superluminal particles and always outside for subluminal ones.
@@
Another relevant insight from our analysis provides the simple discovery of how the creation of subluminal particles, described theoretically [18, 19, 31] and experimentally [37, 32, 38, 39] in stationary subcritical flows, occurs in principle and reconciles with energy conservation. We unveiled that in such set-ups, high-energy particles experience an effective horizon if the flow gets sufficiently close to forming a Killing horizon. For such modes, the Hawking partner, propagating inside the effective horizon, is given by a negative Killing frequency particle, and hence energy conservation will hold as usual.
A final comment must be devoted to the vacuum state. In our analysis, we implicitly assumed a “no drama” scenario for observers crossing the effective horizon or the acoustic/Killing horizon along the “approximant” trajectory. One could interpret this as implicitly assuming the existence of the usual global Unruh vacuum state, similar to what is done in the Bogoliubov method for stationary flows. However, rather than comparing particular states, the tunneling method performs a time-resolved analysis that selects an instantaneous local vacuum to count particles [45]. In other words, as long as a near-horizon vacuum remains valid, the method is predictive. This assumption is justified by the physical intuition of how such a vacuum would be produced in a realistic experiment generating a supersonic flow and is substantiated by experimental data [46].

### v2

@@ VI Particle creation: subcritical flow
Now, we complete our analysis by investigating a subcritical flow for which $|v|<1$ besides $|v_{\mathrm{max}}|\sim O(1)$. Then, no Killing/sonic horizon exists, hence such a flow mimics an horizonless ultra compact object (a quasi-black hole) rather than a black hole, cf. [42, 43] for further references. This is a quite interesting case, given that experimental problems in shallow water waves experiments concerning the stability of horizons led to the prevalent realisation of such “near-critical” flows  [44, 36, 45, 46]. Experiments that nonetheless observed particle production. This led to several numerical [18, 19] and analytical investigations [35] which verified the presence of a particle flux albeit generically characterised by substantial deviations from thermality.
Here we shall re-analyse this phenomenon using the tunneling method with the above introduced approximant trajectory. As we shall see, our technique will be able to capture the spontaneous emission of an analogue ultra compact object in an “Hawking-like” picture, through the tunneling method in a non-relativistic setting like ours. The absence of an acoustic horizon, however, will reflect in a strong non-thermal character of such “pre-Hawking” particle production phenomenon.
It is however important to stress that this effect will not be the only source of particle production for a subcritical flow in the presence of subluminal perturbations. What is described in the following reflects the fact that, for subluminal dispersion, and depending on the asymptotic values of the flow velocity $v(x)$, some high-energy (high $\alpha$) modes enjoy the presence of a turning point, as in figure 1. These modes, despite the absence of a sonic horizon, feel the presence of an EFH, and can be treated the same way as we have done in the supercritical case (because they admit two real solutions inside the EFH and four outside).
However, besides this, there is always a dominant set of low-energy (low $\alpha$) modes, which will not show any turning point, thus admitting four real solutions on the whole $(x,t)$ plane. For such modes, it is possible to show, by analyzing the Bogoliubov coefficients in a $4\times 4$ scattering matrix, that a stimulated processes may occur as well [47]. This latter effect has a completely different root to the one we focus here and could easily be dominant in a noisy experimental environment.
Indeed, this would be more similar in nature to the so called “Klein-paradox” which describes a special case of superradiant amplification for scalar fields impinging on a potential barrier [48]. Such a mechanism requires an incoming wave to start with, while the high-energy modes experiencing an EFH will also probe an effective ergoregion which allows for a spontaneous emission (vacuum instability).
Nonetheless, we feel it is quite interesting to see how a spontaneous channel can be present even in the absence of a full fledged acoustic horizon, as it predicts that even ultra compact horizonless object could be evaporating in non-relativistic matter frameworks (like, for example, the so called Standard Model Extension, see e.g [49]).
Let us then start our analysis with the EFH condition (23).
@@
Before particularising to a specific profile of $v(x)$, we quantify first how far the flow’s most negative value can be from the sonic point, $v=-1$, while still admitting some solution to (23) for our subluminal dispersion relation. Within our setting this yields numerically the upper bound (namely, the solution of (23) for $\alpha=1/2$) [^5] The value for $\alpha$ is in principle arbitrarily chosen. In this setup, the threshold is determined by the value where the dispersion relation becomes multivalued. From a physical point of view, the minimal value represents that basically every object, no matter how far from forming a horizon, would radiate modes $k\sim\Lambda$.
@@
In order to proceed further in our investigation, let us now assume a velocity profile like in figure 8 given by the form
@@
Figure 9: $\kappa(\alpha)$ for different values of $\varepsilon$: increasing $\varepsilon$ reduces the range of $\alpha$ for which (23) has a solution. Moreover, $\kappa(\alpha)$ starts from 00 with a non-vanishing derivative (apart from the critical case $\varepsilon=0$) which implies a deviation from thermality in the emission that will be the topic of the last section
@@ VII Phenomenological Considerations
In this section, we highlight some important aspects and subtleties related to our approach to provide a complete and comprehensive picture. For instance, we analyse the properties of the spectrum and the internal consistency of the methodology.
@@ VII.3 Energy conservation: subcritical flow-subluminal dispersion relation
  • the lower branch of the negative energy mode for the subluminal case (which is the lingering mode in Fig.1) is dispersive around $v=0$ (the origin of the axes in Fig.1). In (16) we have shown that the regular solutions around $v\simeq 0$ consist in two positive-$\Omega$ solutions, one ingoing and one outgoing. However, none of the lower branches of the lingering or the turning mode reach $v=0$, meaning that those branches are non-regular in that region of spacetime. This can be explained by the following fact: for these two branches $|k|$ increases while approaching $v=0$, so much so that for the subluminal dispersion relation $\omega(k)$ becomes imaginary and no real mode of these branches can be further propagated towards $v=0$.
@@ VII.4.2 Subluminal case
In conclusion, we see that quite generically the particle creation process described via the tunelling method and the approximant, is basically insensitive of the near-horizon behaviour which is where the details of the dispersive behaviour of our modes matter most. This resonates with the more complete, but also more convoluted, results of the Bogoliubov approach, where the same robustness to the short-distance details of the process has been found, see cf. [40].
@@ VIII Discussion
This article was devised as a pedagogical primer for Hawking radiation in acoustic gravity, providing an accessible and physically intuitive description of this phenomenon via the tunneling method. Due to its simplicity and versatility, this formalism allowed us to explore several different, albeit sometime idealized, scenarios that may be relevant in analogue gravity set-ups. We investigated three different flow types — supercritical, critical, and subcritical — over which we analysed the particle creation for a scalar field with generalised dispersion relations of subluminal as well as superluminal type. As a result, we found that particle creation can be associated with an effective horizon experienced by a trajectory that interpolates those of the Hawking partners outside and inside the (effective) horizon. Such approximant only fails to track these particles in a region of size below their de Broglie wavelength (where they cannot be treated as separated on-shell particles). The aforementioned effective horizon is always located inside the Killing horizon for superluminal particles and always outside for subluminal ones.
@@
Another relevant insight from our analysis provides the simple discovery of how the creation of subluminal particles, described theoretically [18, 19, 35] and experimentally [44, 36, 45, 46] in stationary subcritical flows, can have a spontaneous component, in addition to a dominant stimulated one. Indeed, we unveiled that in such set-ups, some high-energy modes can experience an effective horizon, if the flow gets sufficiently close to forming a Killing horizon. For such modes, we showed that the Hawking partner, propagating inside the effective horizon, is endowed with a negative Killing frequency, and hence energy conservation can be made to hold as usual in spite of being in vacuum and in a time independent geometry.
A final comment must be made on the vacuum state. In our analysis, we implicitly assumed a “no drama” scenario for observers crossing the effective horizon or the acoustic/Killing horizon along the “approximant” trajectory. One could interpret this as implicitly assuming the existence of the usual global Unruh vacuum state, similar to what is done in the Bogoliubov method for stationary flows. However, rather than comparing particular states, the tunneling method performs a time-resolved analysis that selects an instantaneous local vacuum to count particles [55]. In other words, as long as a near-horizon vacuum remains valid, the method is predictive. This assumption is justified by the physical intuition of how such a vacuum would be produced in a realistic experiment generating a supersonic flow and is substantiated by experimental data [56].
@@ Acknowledgments
Appendix A Tunneling vs. Bogoliubov
This appendix clarifies the relation between the tunneling and the Bogoliubov approach. The main point of our discussion will be to understand why the interpretation of a tunneling rate $\Gamma$ as defined in (13) can be interpreted as a signature of thermality for the observer at infinity ($v(x)=0$) where the particle content is usually determined by evaluating the Bogoliubov coefficients between the near-horizon modes and the modes at infinity. This analysis, being a proof of concept, will be conducted fully in the relativistic case, but we will argue that it can be applied to fields with modified dispersion relations, since the short-range details of the modes will not matter in the evaluation of the Hawking effect. For further details on the derivations, see [57] (chapter 2) and for similar consideration see [31].
Let us then consider a relativistic field $\phi$, solving the massless Klein-Gordon equation in a static, $(1+1)$ dimensional [^9] The very same discussion can be made with a spherically symmetric black hole example , asymptotically flat geometry with a Killing horizon, such as (1),
$\square\phi=0\,.$ (54)
In order to compute the Hawking effect we focus on the outgoing branch of the solutions. This can be described labelling the modes at fixed Killing energy $\Omega$. Both inside and outside the horizon, these modes assume the usual logarithmic non-analyticity in the near-KH limit
$\phi_{\Omega}^{\textsc{kh},\rm out}=\exp\left[-i\Omega t+\Omega\frac{i}{\kappa_{\textsc{kh}}}\ln(x-x_{\textsc{kh}})\right]\,,\quad\phi_{\Omega}^{\textsc{kh},\rm in}=\exp\left[-i\Omega t+\Omega\frac{i}{\kappa_{\textsc{kh}}}\ln(x_{\textsc{kh}}-x)\right]\,,$ (55)
while at infinity, where the geometry is flat, they assume the shape of a plane wave:
$\phi_{\Omega}^{\infty}=\exp\left[-i\Omega(t-x)\right]\,.$ (56)
The Bogoliubov relation between the two (inequivalent) bases $\{\phi_{\Omega}^{\textsc{kh},\rm out}\}$ and $\{\phi_{\Omega}^{\infty}\}$, together with the assumption of the state to be vacuum for the freely-falling observer at horizon crossing, essentially captures the spontaneous particle production of a black hole, that is the Hawking effect.
An intuitive way to compute the Bogoliubov coefficients is to perform a trick à la Unruh [58], where we get rid of the non-analyticity given in (55), building a new set of solutions which is thus analytical across the KH [59]:
$\Phi_{\Omega}^{\pm}=C^{\pm}\left[\phi_{\Omega}^{\textsc{kh},\rm out}\pm e^{\mp\frac{\Omega}{2\kappa_{\textsc{kh}}}}(\phi_{\Omega}^{\textsc{kh},\rm in})^{*}\right]\,.$ (57)
The exponential $e^{\mp\frac{\Omega}{2\kappa_{\textsc{kh}}}}$ that appears in the expression above, represents nothing else than the analytical continuation of the logarithm in (55) across the KH on the upper and lower half complex $x$-plane, depending on the sign of the exponent. The coefficients $C^{\pm}$, instead, are normalization coefficients. The analyticity of $\Phi^{+}_{\Omega}$ ($\Phi^{-}_{\Omega}$) allows us to express this as a sum of only positive (negative) frequency modes at infinity $\phi_{\Omega}^{\infty}$. This means that the vacuum state for the observer at infinity is the same vacuum defined through the $\Phi^{+}_{\Omega}$. So, the evaluation of the Bogoliubov coefficients can be equivalently made, instead of between $\{\phi_{\Omega}^{\textsc{kh},\rm out}\}$ and $\{\phi_{\Omega}^{\infty}\}$, just between $\{\phi_{\Omega}^{\textsc{kh},\rm out}\}$ and $\{\Phi_{\Omega}^{\pm}\}$.
As a matter of fact, the mean value of the number operator turns out to be given by [57]:
$-C^{-}=-\langle\Phi^{-}_{\Omega},\phi_{\Omega}^{{\textsc{kh}},\rm out}\rangle=\int\frac{{\rm d}\bar{\Omega}}{2\pi}|\beta_{\Omega\bar{\Omega}}|^{2}=\langle\hat{N}_{\Omega}\rangle\,.$ (58)
Taking into account the normalization condition of $\Phi^{\pm}_{\Omega}$ we get:
$\langle\hat{N}_{\Omega}\rangle=\int\frac{{\rm d}\bar{\Omega}}{2\pi}|\beta_{\Omega\bar{\Omega}}|^{2}=\frac{1}{e^{2\pi\Omega/\kappa_{{\textsc{kh}}}}-1}\,,$ (59)
which is a Bose-Einstein distribution, underlying a thermal character of the particle number $\hat{N}_{\Omega}$ (the expetation value is taken on the freely falling vacuum) with temperature $T=\kappa_{{\textsc{kh}}}/2\pi$, that is the Hawking temperature. However, this is not the distribution which an observer far from the horizon would detect since, due to the curved geometry, the propagation affects the particle content in an energy-dependent way through the so-called grey-body factor $\Gamma_{\rm grey}(\Omega)$ that quantifying the probability of a single particle of energy $\Omega$ to overcome the gravitational potential and eventually reach infinity. The resulting spectrum is given by [59]:
$\frac{\Gamma_{\rm grey}(\Omega)}{e^{2\pi\Omega/\kappa_{{\textsc{kh}}}}-1}\,.$ (60)
Let us underline that the grey-body factor always affects the propagation of the outgoing modes, apart from the particular case of a massless relativistic particle in $(1+1)$-dimensional geometry, because any such metric is conformally flat [14].
The equivalence between the tunneling and the Bogoliubov approach relies on the fact that the source of thermality is mathematically given by the non-analyticity of the outgoing modes near the horizon. The tunneling rate (13) represents exactly the computation of this factor, since it involves the imaginary part of the phase of $\phi$. This equivalence is possible due to the exact WKB character of the solution near the horizon. In other words $\phi_{\Omega}^{{\textsc{kh}},\rm out}$, takes the form given in (55) without relying on any approximation, due to the exponential peeling.
Usually, the tunneling rate $\Gamma$ has a statistical interpretation [60]. In this case, one can interpret the existence of an outgoing particle as the sum of two processes
$\phi_{\Omega}^{{\textsc{kh}},\rm out}=C^{-}\Phi^{-}_{\Omega}+C^{+}\Phi^{+}_{\Omega}\,,$ (61)
which tells us that a particle $\phi_{\Omega}^{{\textsc{kh}},\rm out}$ can be originated by a the outgoing flux with probability $-C^{-}=P_{\rm em}$ and by the reverse process with probability $C^{+}=P_{\rm abs}$. The ratio between those two processes gives us exactly
$\Gamma=\frac{P_{\rm em}}{P_{\rm abs}}=\frac{\langle\phi_{\Omega}^{{\textsc{kh}},\rm out},\Phi^{-}_{\Omega}\rangle}{\langle\phi_{\Omega}^{{\textsc{kh}},\rm out},\Phi^{+}_{\Omega}\rangle}=e^{-2\pi\Omega/\kappa_{\rm KH}}\,.$ (62)
From the formula above we recognize immediately that the numerator and the denominator of the ratio exactly correspond to the ratio of the $\alpha$ and $\beta$ Bogoliubov coefficients:
$\Gamma=\frac{\int\frac{{\rm d}\bar{\Omega}}{2\pi}|\beta_{\Omega\bar{\Omega}}|^{2}}{\int\frac{{\rm d}\bar{\Omega}}{2\pi}|\alpha_{\Omega\bar{\Omega}}|^{2}}=e^{-2\pi\Omega/\kappa_{\rm KH}}\,.$ (63)
This, together with the completeness relation:
$1=\int\frac{{\rm d}\bar{\Omega}}{2\pi}\,(|\alpha_{\Omega\bar{\Omega}}|^{2}-|\beta_{\Omega\bar{\Omega}}|^{2})$ (64)
allows us to derive
$\int\frac{{\rm d}\bar{\Omega}}{2\pi}|\beta_{\Omega\bar{\Omega}}|^{2}=\frac{1}{e^{2\pi\Omega/\kappa_{{\textsc{kh}}}}-1}$ (65)
as obtained above.
This confirms that, through the tunneling rate, one can directly read the relation between the non-analyticity of the modes and the thermal character of the state off. If $\Gamma$ turns out to be a Boltzmann factor, one finds the temperature automatically. Morevoer, this analysis helps to understand where the choice of vacuum plays a role in a context like the tunneling formalism. In fact, such a Boltzmann factor gives rise to a thermal state in the analytical basis only when the non-analytical one is set in vacuum and the number of particle can be identified as the integral $|\beta|^{2}$ only. If the state is not taken to be vacuum for the freely falling observer, this is not true anymore, and, despite $\Gamma$ assumes the same shape, the shape of the state at infinity cannot be directly inferred from it in the same way.
Additionally, let us emphasize that this calculation can be performed locally near the horizon. However, as we have already pointed out, this does not correspond exactly to the detection at infinity, due to propagation effects. This is of particular importance when we consider modified dispersion relations where, even for the massless case, one expects a nontrivial effect from the propagation. Therefore, the results of this article cannot be quantitatively compared, beyond the zeroth order in $\alpha$, with the Bogoliubov coefficient approach, without a previous estimation of the effect of the modes’ propagation up to infinity.